OpenStudy (anonymous):

Choose the ordered pair that is the solution to the system of equations. x + y = 4 y = 2x + 1

OpenStudy (haseeb96):

x+y=4 ...........eq1 y=2x+1.............eq2 now put eq2 into eq1 x+2x+1 =4 3x=4-1 3x=3 x=1 put x=1 into eq2 y=2(1)+1 y=3 solution set:- (1,3) this is the correct answer